OpenStudy (anonymous):

Square root of 3d^2-8= d-2

OpenStudy (akshay_budhkar):

\[\sqrt{3d^2-8}=d-2\]

OpenStudy (akshay_budhkar):

square it 3d^2-8=d^2-4d+4....

OpenStudy (akshay_budhkar):

2d^2+4d-12=0

OpenStudy (akshay_budhkar):

d^2+2d-6=0

OpenStudy (anonymous):

My options are 3/4 , 3/-4 , 3 , 4 .

OpenStudy (akshay_budhkar):

is none of thse a option?

OpenStudy (anonymous):

N0 .

OpenStudy (akshay_budhkar):

wait a min

OpenStudy (akshay_budhkar):

i believe the question is wrong you will get the answer in root

OpenStudy (akshay_budhkar):

plz comfirm the question... if it is it then i must say it is wrong.. sorry:(

OpenStudy (anonymous):

what is this?

OpenStudy (akshay_budhkar):

have i done a mistake?????? its not matching :(

OpenStudy (anonymous):

no solution to this one

OpenStudy (akshay_budhkar):

ok so d ques is wrong... i was worried..sorry guys.. phew

OpenStudy (anonymous):

the line \[y=x-2\] lives below the curve \[y=\sqrt{3x^2-8}\]

OpenStudy (anonymous):

I got it . The answer is 3/4 .

OpenStudy (akshay_budhkar):

Lyssah621how??????????

OpenStudy (anonymous):

oh no it is not

OpenStudy (anonymous):

3/4 is not in the domain of \[\sqrt{3x^2-8}\]